OpenStudy (anonymous):

A jar contains jelly beans that are red, yellow, blue, and purple. Several people reach in and grab handfuls of jelly beans, then collectively tally the number of each color: 15 red, 20 yellow, 12 blue, and 8 purple. If someone pulls out 22 jelly beans, about how many would you expect to be yellow? A. 8 B. 13 C. 20 D. 4

OpenStudy (anonymous):

step 1: Add up your reds,yellows, blues, and purples step 2: you create an equation -->\[\frac{ 22 }{ 52}\]= \[\frac{ x }{ 20 }\] step3: you cross multiply --> 52x = 440 step 4: divide both sides by 52 step 5: your answer is A

OpenStudy (anonymous):

Thanks:)

OpenStudy (anonymous):

The fraction yellow is approximately f=20/55 Withdrawing 22 gives expected number (f)(22)= 22(20)/55 = 8

OpenStudy (anonymous):

so @chelseybell96 is correct.